OpenStudy (anonymous):

Find an equation for the line tangent to the graph of: f(x)=6xe^x at the point (a,f(a)) for a =2 y= ?

OpenStudy (anonymous):

did you take the derivative?

OpenStudy (anonymous):

I was doing implicit differentiation

OpenStudy (anonymous):

product rule you mean?

OpenStudy (anonymous):

implicit differentiation is just for when your function is a function of x and also y

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

So i take the derivative?

OpenStudy (anonymous):

yes, the derivative tells you the slope of the tangent line (for any x)

OpenStudy (anonymous):

tell me what you get for the derivative...

OpenStudy (anonymous):

f(x)'=6xe^x=6e^x

OpenStudy (anonymous):

good

OpenStudy (anonymous):

oops that was a plus

OpenStudy (anonymous):

I figured:)

OpenStudy (anonymous):

f(x)'=6xe^x+6e^x

OpenStudy (anonymous):

so we want the slope at x=2

OpenStudy (anonymous):

18e^2

OpenStudy (anonymous):

is m the slope use in y=mx+b with x=2 and y = 6*(2)*e^2 to find b

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

you found the slope of any tangent line: 6xe^x+6e^x the slope of the tangent line for x=2 is: 6*2*e^2 + 6*e^2

OpenStudy (anonymous):

18e^2

OpenStudy (anonymous):

SO that is the slope when x=2

OpenStudy (anonymous):

yep;)

OpenStudy (anonymous):

Ohh haha

OpenStudy (anonymous):

we can use that to find the equation of the tangent line that goes through that point...

OpenStudy (anonymous):

y = 18*e^2 *x +b

OpenStudy (anonymous):

and the point we're looking at is (2, 6*2*e^2)

OpenStudy (anonymous):

sub.s those points in for x and y, and solve for b:)

OpenStudy (anonymous):

OKay cool thank you

OpenStudy (anonymous):

ok? all good? any questions?

OpenStudy (anonymous):

Sure!